34. A uniform ladder of mass m rests against a frictionless vertical wall at an angle of 60°. The lower endth mass M=2m attempts to climb the ladder. What fraction of the length L of the ladder will
the student have reached when the ladder begins to slip?
1
Expert's answer
2020-08-18T10:08:35-0400

"H/V=0.4\to H=0.4V"


"H=F"


"V-mg-2mg=0\to V=3mg"


"H=0.4\cdot3mg=1.2mg"


"-mg(L/2)\cos60°-2mgx\cos60°+1.2mgL\sin60°=0"



"-(L/2)\cos60°-2x\cos60°+1.2L\sin60°=0"


"-L/4-x+1.039L=0\to\frac{x}{L}=0.789"


So,


"x\approx79\%" of "L".